OpenStudy (anonymous):

Need help with linear approximation... f(x) = x^1/3, find equation of tangent line when x = 27.

OpenStudy (anonymous):

take the derivative it is \[f'(x)=\frac{1}{3}x^{-\frac{2}{3}}=\frac{1}{3\sqrt[3]{x^2}}\]

OpenStudy (anonymous):

then for the point on the graph find \[f(27)=\sqrt[3]{27}=3\] so the point is (27,3) and to find the slope find \[f'(27)=\frac{1}{3\sqrt[3]{27^2}}=\frac{1}{3\times 3^2}=\frac{1}{27}\]

OpenStudy (anonymous):

then use the point slope formula to find the equation for the line

OpenStudy (anonymous):

can you show me that part as well? I'm kinda lost in this part as well.

OpenStudy (anonymous):

use \[y-y_1=m(x-x_1)\] with \[m=\frac{1}{27}, x_1=27,y_1=3\] a direct substitution will give you the equation for the line

OpenStudy (anonymous):

alright, thx a bunch